OpenStudy (dumbsearch2):

May you verify if my answer is correct ? An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the object's maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t2 + 64t + 80? A. 272 B. 144 C. 1232 D. 616

OpenStudy (dumbsearch2):

I got B. 144 by plugging in or t, however I'm really not sure.

OpenStudy (dumbsearch2):

*for

OpenStudy (aum):

h(t) = -16t^22 + 64t + 80 is the equation of a parabola. To find the maximum, you need to find the vertex of the parabola. The x-coordinate of the vertex occurs at -b / 2a So it occurs when t = -64 / (2*(-16)) = -64 / -32 = 2 sec. put t = 2 in h(t) to find the max height. h(2) = -16(2)^2 + 64(2) + 80 = -64 + 128 + 80 = 144 feet.

OpenStudy (campbell_st):

correct